# 3D ISM structure challenges the Serkowski relation

### v1

@@ 3.1 Polarimetric data selection – Statistical sample
A recent agglomeration of multi-band optical polarization data (Panopoulou et al., 2023), which is the largest to date, allows us, for the first time, to study in a statistical fashion the impact of the 3D ISM structure of the Galaxy on the Serkowski relation, and the consequences for magnetic field and dust physics studies. In order to fit multiple-cloud models to polarization data it is optimal to use polarization measurements across a broad optical wavelength range. However, for most stars in the literature, measurements exist only in a single or in few bands. Even so, it is possible to evaluate the validity of the Serkowski relation in a statistical manner, by studying an ensemble of sources. We designed such an experiment to test whether stars in single-cloud LOSs are better fit by the Serkowski relation than stars in multiple-cloud LOSs. For acquisition of archival optical polarization measurements for the statistical comparison, we explored Table 6 of the compiled polarization catalog of Panopoulou et al. (2023), which provides polarimetry and distances from Gaia EDR3 (Bailer-Jones et al., 2021) for $\sim 42,000$ sources. We excluded stars with distance greater than 1.25 kpc, as this is the limit of our dust map (see Sect. 3.2). As distance, we take the average of the values (‘r_med_geo’,‘r_med_photogeo’) provided in the catalog. We excluded measurements in filters with no well-defined effective wavelength ($\lambda_{eff}$), therefore, the excluded filters are “0”, “20”, “23”, “51” which correspond to “No filter or unclear”, “weighted-mean”, 0.735-0.804 $\mu m$, 0.5-0.85 $\mu m$, respectively in the catalog. For targets with multiple measurements in the same filter we used the weighted mean and standard deviation of the $q$ measurements, according to
$q_{\text{mean}}=\frac{\sum(q_{i}\cdot w_{i})}{\sum w_{i}}\,\,and$ (4)
@@
with $w_{i}=1/\sigma_{qi}^{2}$, and $\sigma_{qi}$ the uncertainty of the measurement in $q$, and similarly for $u$. We excluded measurements with debiased $P/\sigma_{P}<2$. Additionally, we did not consider measurements with polarization uncertainty $\sigma_{P}<0.1\%$ because they are likely underestimated 333probably they refer exclusively to the statistical uncertainties, ignoring the systematic ones, as the most reliable polarization standards to date (Blinov et al., 2023) are only accurate down to $0.1\%$. After all the cuts, we selected targets that have surviving measurements in at least three bands. We note that we did not impose any cut for intrinsically polarized stars (although the Panopoulou et al. (2023) catalog provides this information), as these targets may have been flagged as intrinsically polarized simply because they do not follow the Serkowski relation. However, the final sample does not contain any star that is flagged as intrinsically polarized. After all the cuts, the final sample contains 223 sources with measurements in at least three optical bands.
@@ 3.2 Extinction
We collected extinction data for our target stars from the Edenhofer et al. (2024) dust map. We choose this map, as it provides good angular resolution (14’), as well as, parsec-scale distance resolution. The map is available in two versions, one that extends out to 1.25 kpc from the Sun, and one that extends up to 2 kpc, but uses less data for the dust profile reconstruction. We used the 1.25 kpc version, as it should be more accurate, and we did not want to introduce any additional sources of noise. Extinction values in the Edenhofer et al. (2024) map are expressed in the units used by Zhang et al. (2023), the work upon which the dust map was based. We did not renormalize the extinction values in another units system, as we are not interested in the absolute values, but in variations in the dust distribution along the LOS.
@@
For each of our targets, we examined the cumulative extinction ($E$) vs distance ($D$) profiles by eye, in order to identify the number of significant steps ($N_{C}$) in the profile, which presumably correspond to distinct dust clouds. For targets where more than 3 clouds were identified, or the dust profile was very complex, we set $N_{C}>3$. The number of target stars in each subgroup with different number of identified clouds along the LOS are #$\#(N_{C}=1)=49$, #$\#(N_{C}=2)=74$, #$\#(N_{C}=3)=70$, #$\#(N_{C}>3)=30$. The targets that correspond to each $N_{C}$ are distributed over different positions on the sky (Fig. 7).
@@ 3.3 Fitting the Serkowski relation
In order to evaluate the performance of the Serkowski relation in our statistical sample, we used an extremely powerful, if unconventional in Serkowski studies, technique: we fit our data in the Stokes parameters space $q-u$, exploiting its power to simultaneously trace changes in the profiles of both the degree of polarization and EVPA. Traditionally, the EVPA is not taken into account while fitting the Serkowski formula, thus part of the information imprinted in the polarization signal is ignored. Although the variability of the EVPA with wavelength has been hypothesized to be connected to the existence of multiple clouds (e.g. McMillan & Tapia, 1977; Whittet, 2015), $q-u$ fitting has never been employed to trace the various components (clouds) contributing to the $P-\lambda$ and $\theta-\lambda$ profiles.
As a first step we converted Eq. 1 to the $q-u$ space, with the introduction of $\theta$ in the model as an additional free parameter. Therefore, our model becomes
@@
where $N_{\lambda}$ the number of different bands where we had measurements for each target, $q_{i}$, $u_{i}$ the measurement in a given band, $\hat{q_{i}}$, $\hat{u_{i}}$ our model as introduced in Eq. 6a, 6b, and $\sigma_{q_{i}}$, $\sigma_{u_{i}}$ the uncertainties of the measurements in $q$ and $u$ respectively. The terms $L_{q}$ and $L_{u}$ are the reduced $\chi^{2}$ metric for $q$ and $u$ respectively. The terms $L_{q}/L_{u}$ and $L_{u}/L_{q}$ were introduced to ensure a balanced contribution from $L_{q}$ and $L_{u}$. For example, if the objective was to minimize just $L_{q}+L_{u}$, the optimal value $V$ could be achieved with $L_{q}=V$ and $L_{u}=0$, $L_{u}=V$ and $L_{q}=0$, or any combination in between. By introducing $L_{q}/L_{u}$ and $L_{u}/L_{q}$, we penalize the model when it favors solutions where $L_{q}=0$ or $L_{u}=0$, thus promoting a more balanced fit.
For the modelling procedure we used the Optuna package in python (Akiba et al., 2019). Optuna is an automatic hyperparameter optimization software framework, originally developed for machine learning hyperparameter-tuning purposes. However, it can also address general optimization problems through the use of a Bayesian optimization framework. We found it highly effective in multi-parameter optimization with few data points and many parameters, such as the case of the $q-u$ modelling, and more efficient in finding the best solution than traditional algorithms such as Markov-Chain Monte-Carlo (MCMC). We used the default options of Optuna with n_trials = 5000 for our fits. Moreover, we adopted the following priors. For $P_{max}$ we initially assumed a rather loose Gaussian prior distribution with a mean of $3\%$ and standard deviation of $3\%$, and enforced it to be positive. For $\lambda_{max}$ we assumed a gaussian prior distribution with the mean 0.6 $\mu m$ and standard deviation 0.2 $\mu m$, and ensured it was always positive. For $K$ we imposed a Gaussian prior distribution with mean $1.68\cdot\lambda_{max}$ and standard deviation $0.13\cdot\lambda_{max}$ (Wilking et al., 1980), and ensured it was always positive. For $\theta$ we assumed a Gaussian prior distribution with mean the average of the polarization values of measurements in individual bands (derived from the average $q$ and $u$ as in Eq. 4 and converted to $\theta$ using Eq. 3b) and $10^{o}$ standard deviation .
@@ 3.4 Fitting multiple clouds
In order to fit multiple clouds (as in the examples of Sect. 4.1) we used again the $q-u$ fitting technique, following the same logic discussed in Sect. 3.3. The multiple cloud model is derived from Equations 2 and 6, by adding up the $q$ and $u$ components of the individual clouds.
When trying to determine the correct parameters for a model with two or more clouds, the number of parameters is increased (four parameters per cloud), resulting in a substantial number of possible combinations and a vast parameter space to explore. In such situations, identifying the optimal combination of parameters that minimize the cost function can be challenging. To address this, we employed the following procedure. Initially, instead of directly using Optuna, we utilized an MCMC algorithm to rapidly explore the parameter space. Given that the two clouds are interchangeable, we imposed the condition $\lambda_{max_{1}}>\lambda_{max_{2}}$ to focus the algorithm on a single solution. Subsequently, we visually examined the corner plots of the parameters generated by the MCMC algorithm to establish reasonable constraints for the Optuna algorithm, thereby significantly narrowing the search space. Finally, we ran the fit using Optuna, imposing Gaussian priors on the parameters, based on the mean and the spread of the parameters identified by the MCMC algorithm. We chose Optuna to derive the final set of parameters for the model, as it directly provides the optimal set of parameters by minimizing Eq. 7. The fits are evaluated with $\chi^{2}_{\nu}$ (Eq. 9).
@@ 4.1 Two revealing examples
To demonstrate observationally the impact of the LOS variations of $\lambda_{max}$ and EVPA on the Serkowski relation, we employed archival spectropolarimetric data (Bagnulo et al., 2017). We identified two revealing cases, where the multi-wavelength polarization data show clear evidence of the LOS effects.
The LOS toward the main sequence star HD93222 with spectral type O7 represents a prominent example of what the $P-\lambda$, and $\theta-\lambda$ profiles look like when 3D effects are significant. Firstly, we observe that the $P-\lambda$ deviates from the typical Serkowski curve (Fig. 2 middle panel). Fitting the $P-\lambda$ data with the Serkowski relation requires unphysical values for both $\lambda_{max}$, and $K$ (Table 1), which still yield a merely adequate fit. In addition, $\theta$ varies smoothly with wavelength $\lambda$, which is unexpected when the polarization is dominated by a single cloud with a well defined mean magnetic field. Thus, both $P$ and EVPA show an atypical behaviour with $\lambda$ for this star. Using 3D dust extinction maps (Edenhofer et al., 2024), we verified the existence of multiple clouds along the LOS toward the HD93222 star. The clouds are identified as abrupt increments in extinction as a function of distance (left panel in Fig. 2). The average extinction profile shows three prominent clouds located at around 0.25, 1.35 and 1.8 kpc. The target star, whose estimated distance is 2.44 kpc (Bailer-Jones et al., 2021), lies beyond these clouds, hence its polarization signal carries information from all of them. We fit the data with models with different number of clouds. Although the extinction profile exhibits three distinct steps, the data can be described equally well with both a 2- and a 3-cloud model. This may indicate that two out of three clouds have similar properties. We proceed with the simplest case of a 2-cloud model.
@@
Figure 2: Example of a case where both the $P-\lambda$ and the $\theta-\lambda$ profiles deviate from Serkowski expectations. We are able to fit both profiles by introducing a 2-Cloud model. Data from Bagnulo et al. (2017) for HD93222. Left: Extinction vs distance profile taken from the 2kpc version of the map by Edenhofer et al. (2024). The vertical red line corresponds to the distance of the source. Middle: Data in the $P-\lambda$ space, together with three fitted models: Solid black line: Fit using Eq. 1 in $P-\lambda$ space only. Dashed blue line: Best-fit 1-Cloud model in the $q-u$ space (Eq. 6). Dashed-dotted red line: Best-fit 2-Cloud model in the $q-u$ space (Eq. 6). Dotted orange lines correspond to the two individual components of the 2-Cloud model. Right: Data in $\theta-\lambda$ space together with the three models and the individual components of the 2-Cloud model. Parameters of the fits are quoted in Table 1.
@@
As a second example, we show results for the star HD163181, which is an intermediate-size luminous supergiant of spectral class O9.5 (Fig. 3). In this case, the observed degree of polarization follows the Serkowski formula, but the EVPA varies with wavelength. The 3D extinction profile has two significant steps, hence indicating the existence of more than one dominant clouds along this LOS. Assuming the standard Serkowski curve with a single dominant cloud can yield good fits to the data in the $P-\lambda$ space (Bagnulo et al., 2017), but it fails to explain the observed variability of the EVPA; we verified this by fitting a one cloud model in the $q-u$ plane (dashed blue line in Fig. 3). However, when we considered two clouds along this LOS, as the 3D extinction map suggest, we were able to capture the observed variability of both $P$ and $\theta$ with wavelength (red dashed dotted line in Fig. 3 middle panel).
@@
Figure 3: Example of a case where the $P-\lambda$ profile follows the Serkowski relation, but the EVPA does not. By employing a model with two components in the LOS, it is possible to fit both $P$ and $\theta$. Data from Bagnulo et al. (2017) for HD163181. Panels and lines similar to Fig. 2. Parameters of the fits are quoted in Table 1
4.2 Statistical sample
We grouped the sample of sources found in Sect. 3.1 into subsamples depending on the number of clouds ($N_{C}$) lying in their LOS, as derived from the 3D extinction map. We obtained four main subsamples: targets in LOSs with one dominant cloud, $N_{C}=1$; with two clouds, $N_{C}=2$; three clouds $N_{C}=3$; and with more than three clouds, $N_{C}>3$. We evaluated the performance of the Serkowski relation in all of the targets as described in Sect. 3.3. We constructed the cumulative density functions (CDFs) of $\chi^{2}_{\nu}$ for the different subsamples (Fig. 4), and we applied the Kolmogorov-Smirnov (KS) test to identify any statistical discrepancies between the $\chi^{2}_{\nu}$ distributions of the $N_{C}=1$ subset and the other three subsets.
Figure 4: CDFs of $\chi^{2}_{\nu}$ for different subsamples as described in each label.
We found that the $\chi^{2}_{\nu}$ distribution of the $N_{C}=1$ subset is significantly different from those of subsets with $N_{C}=2$ or $N_{C}=3$. On the contrary, the $\chi^{2}_{\nu}$ distributions of the $N_{C}=1$ subset and of the $N_{C}>3$ subset are not discrepant. The $p-$values of the KS tests for all comparisons are quoted in Table 2, along with the median $\chi^{2}_{\nu}$ for the fits of each subsample. Despite the sparse nature of the data (measurements in only $\sim 3.5$ different filters on average per target) the sample is sufficiently large to allow the statistical distinction of LOSs with two or three clouds from those with only a single cloud. The deterioration of the Serkowski fits in the multi-cloud LOSs is mainly driven by the behaviour of the EVPA with wavelength, and is made clear thanks to our $q-u$ space fits. In contrast, if we perform the fits only on $P-\lambda$, it is not possible to discriminate between LOSs with and without multiple clouds with current data. When the structure in the LOS is too complicated ($N_{C}>3$), there is no statistical difference in the polarization profiles compared to the single cloud case. This is not surprising, as with an increasing number of clouds the variations induced by individual screens tend to average out. Even deviations in the EVPA, which should be more prominent, would become smeared and difficult to detect, especially with limited data points across the spectrum.
Table 2: $p$ values of the KS test between the subsample of targets with $N_{C}=1$ and other subsamples. See Sect. 4.2 for details.
Subsamples $p$-value Median $\chi^{2}_{\nu}$
$N_{C}=1$ 1 1.20
$N_{C}=2$ 1.5 $\cdot 10^{-3}$ 2.59
$N_{C}=3$ 2.1 $\cdot 10^{-2}$ 2.17
$N_{C}=2$ or 3 $1.8\cdot 10^{-3}$ 2.51
$N_{C}>1$ 2.1 $\cdot 10^{-2}$ 1.97
$N_{C}>3$ $1.5\cdot 10^{-1}$ 0.86
4.3 On the parameters of the Serkowski formula

### v3

@@ 3.1 Polarimetric data selection – Statistical sample
A recent agglomeration of multiband optical polarization data (Panopoulou et al., 2025), which is the largest to date, allows us, for the first time, to study in a statistical fashion the impact of the 3D ISM structure of the Galaxy on the Serkowski relation, and the consequences for magnetic field and dust physics studies. In order to fit multiple-cloud models to polarization data, it is optimal to use polarization measurements across a broad optical wavelength range. However, for most stars in the literature, measurements exist only in a single band or a few bands. Even so, it is possible to evaluate the validity of the Serkowski relation in a statistical manner, by studying an ensemble of sources. We designed such an experiment to test whether stars in single-cloud LOSs are better fit by the Serkowski relation than stars in multiple-cloud LOSs.
To acquire archival optical polarization measurements for the statistical comparison, we explored Tables 4 and 6 of the compiled polarization catalog of Panopoulou et al. (2025), which provide polarimetry and distances from Gaia EDR3 (Bailer-Jones et al., 2021), respectively, for $\sim 42,000$ sources. We excluded stars with distances greater than 1.25 kpc, as this is the limit of our dust map (see Sect. 3.2). As a distance, we took the average of the values (‘r_med_geo’,‘r_med_photogeo’) provided in the catalog. We excluded measurements in filters with no well-defined effective wavelength ($\lambda_{eff}$); therefore, the excluded filters are “0,” “20,” “23,” and “51,” which correspond to “No filter or unclear,” “weighted-mean,” 0.735-0.804 $\mu m$, and 0.5-0.85 $\mu m$, respectively, in the catalog. For targets with multiple measurements in the same filter, we used the weighted mean and standard deviation of the $q$ measurements, according to
$q_{\text{mean}}=\frac{\sum(q_{i}\cdot w_{i})}{\sum w_{i}}\,\,$ (4)
@@
with $w_{i}=1/\sigma_{qi}^{2}$, and $\sigma_{qi}$ the uncertainty of the measurement in $q$, and similarly for $u$. We excluded measurements with debiased $P/\sigma_{P}<2$. Additionally, we did not consider measurements with a polarization uncertainty of $\sigma_{P}<0.1\%$ because they are likely underestimated. After all the cuts, we selected targets that have surviving measurements in at least three bands. We note that we did not impose any cut for intrinsically polarized stars (although the Panopoulou et al. (2025) catalog provides this information), as these targets may have been flagged as intrinsically polarized simply because they do not follow the Serkowski relation. However, the final sample does not contain any star that is flagged as intrinsically polarized. The final sample contains 223 sources with measurements in at least three optical bands.
The reasoning for excluding data with $\sigma_{P}<0.1\%$ is presented below. Initially, we considered not imposing any restrictions on the level of $\sigma_{P}$. However, we found that error bars below $\sim 0.1\%$ are likely underestimated on average. If the error bars accurately represented the uncertainties in the measurements, the $\chi^{2}_{\nu}$ would remain independent of the error bar level when analyzing targets within the same subsample (i.e., those behind the same number of clouds). To test this hypothesis, we examined the targets in the subgroup with $N_{C}=1$. We chose this particular subsample, as we know not only that the $\chi^{2}_{\nu}$ should be flat with respect to the error bars, but also that it should be $\sim 1$ for targets behind a single cloud. We fit the model described in Sect. 4.1 and, for each target, plotted the average error bar value across different filters against the corresponding $\chi^{2}_{\nu}$. We demonstrate in the left panel of Fig. 2 that targets with smaller error bars exhibit higher $\chi^{2}_{\nu}$, revealing a clear, nonconstant correlation between the two quantities. $\chi^{2}_{\nu}$ seems to approach the expected value of $\sim 1$ for $\sim\sigma_{P}\geq 0.1\%$. Based on the above findings, we gradually excluded points with lower reported uncertainties and repeated the fit described in Sect. 4.1, until the expected behaviors were satisfied. We found that the dependence practically disappears and $\chi^{2}_{\nu}$ approaches the expected value of $\sim 1$ for targets with average $\sigma_{P}\geq 0.1\%$ (Fig. 2, right). Therefore, we discarded all measurements with $\sigma_{P}<0.1\%$.
Figure 2: Left: Average reported uncertainties in polarization, $\overline{\sigma_{P}}$, versus $\chi^{2}_{\nu}$ for targets that survive the cuts described in Sect. 3.1, but without any restriction for the error bars. The black line is a linear fit for all the data in log-log space. The dashed blue line is similar, but excluding the points in the top left of the plot. Right: Similar to left, but with the restriction for the error bars of $\sigma_{P}\geq 0.1\%$.
@@ 3.2 Extinction
We collected extinction data for our target stars from the Edenhofer et al. (2024) dust map. We chose this map, as it provides a good angular resolution (14’) as well as a parsec-scale distance resolution. The map is available in two versions, one that extends out to 1.25 kpc from the Sun and one that extends up to 2 kpc, but uses less data for the dust profile reconstruction. We used the 1.25 kpc version, as it should be more accurate, and we did not want to introduce any additional sources of noise. Extinction values in the Edenhofer et al. (2024) map are expressed in the units used by Zhang et al. (2023), the work upon which the dust map was based. We converted the extinction values to represent the extinction in the $V$ band ($A_{V}$) based on the conversion table provided by Zhang et al. (2023).
@@
For each of our targets, we examined the “mean” (Edenhofer et al., 2024) cumulative and differential $A_{V}$ versus distance ($D$) profiles by eye, in order to identify the number of significant steps in the profile, which presumably correspond to distinct dust clouds ($N_{C}$). For targets for which more than three clouds were identified, or the dust profile was very complex, we set $N_{C}>3$. The number of target stars in each subgroup with a different number of identified clouds along the LOS are #$\#(N_{C}=1)=49$, #$\#(N_{C}=2)=74$, #$\#(N_{C}=3)=70$, #$\#(N_{C}>3)=30$. The targets that correspond to each $N_{C}$ are distributed over different positions on the sky (Fig. 3).
We considered the possibility of using a more objective metric to characterize the number of dust clouds along each LOS, as each step in the dust profiles is typically not equally prominent, and in many cases numerous small steps contribute to the cumulative extinction. Therefore, we explored a method similar to the one used by Panopoulou & Lenz (2020), who identified the number of clouds in Hi data by detecting peaks in the differential Hi column density profile and then normalizing each peak by the height of the tallest peak among them. In our case, the normalized number of clouds in each LOS would be $N_{C}^{norm}=\sum\frac{A_{V}^{i}}{A_{V}^{max}}$, with $A_{V}^{i}$ the extinction of a given peak in the LOS, and $A_{V}^{max}$ the extinction of the highest peak in the differential $A_{V}$ versus $D$ profile. However, this was not useful in our study for the following reasons. 1) $N_{C}^{norm}$ is ambiguous. For example, $N_{C}^{norm}=2$ could signify two equally prominent peaks, or one prominent peak plus two peaks half its height, or even one prominent peak plus three peaks a third of its height, and so on. This lack of clarity reduces its utility. 2) The profiles in the Edenhofer et al. (2024) map are sometimes irregular, with many minor peaks that may not represent real features but instead result from sampling noise. Additionally, peaks can appear very close together, making it difficult to distinguish between separate clouds and variations within the same cloud. Although we considered smoothing the profiles to address these issues, we found that the number and relative height of the resulting peaks were highly sensitive to the parameters chosen for smoothing, introducing further ambiguity. 3) Even when one peak appears higher than another, this does not necessarily correlate with increased polarization. The dust-induced polarization of aligned grains depends largely on factors beyond peak height, such as dust density, grain alignment efficiency, magnetic field coherence, and the inclination of the magnetic field with the LOS.
In our attempts to devise an objective method of peak counting, we found that the resulting numbers were often ambiguous or lacked physical meaning. Consequently, we opted for a visual approach. Although this method has its uncertainties, it is self-consistent and allows statistical patterns to emerge in a dataset as large as ours. Figure 4 illustrates examples of dust profiles that we assigned to different subsamples.
Figure 3: Positions of targets depending on their $N_{C}$. Maps are in Galactic coordinates, in mollweide projection, centered on (0,0). Parallels are drawn every 30o. Meridians are drawn every 60o. Longitude increases toward the right-hand side. Colors correspond to the average $\lambda_{max}$ of each area. Areas marked with “A” through “G” contain stars that are located behind a single cloud and multiple clouds. See discussion in Sect. 4.3.
Figure 4: Dust profiles for four example targets belonging in four different subsamples. From left to right: $N_{C}=1$, $N_{C}=2$, $N_{C}=3$, $N_{C}>3$. Top row: Cumulative $A_{V}$ vs. $D$. Bottom row: Differential $A_{V}$ vs. $D$.
@@ 3.3 Fitting the Serkowski relation
In order to evaluate the performance of the Serkowski relation in our statistical sample, we used an extremely powerful, if unconventional in Serkowski studies, technique: we fit our data in the Stokes parameters space $q-u$, exploiting its power to simultaneously trace changes in the profiles of both the degree of polarization and EVPA. Traditionally, the EVPA is not taken into account while fitting the Serkowski formula, and thus part of the information imprinted in the polarization signal is ignored. Although the variability of the EVPA with wavelength has been hypothesized to be connected to the existence of multiple clouds (e.g. McMillan & Tapia, 1977; Whittet, 2015), $q-u$ fitting has never been employed to trace the various components (clouds) contributing to the $P-\lambda$ and $\theta-\lambda$ profiles.
As a first step, we converted Eq. 1 to the $q-u$ space. This step allows us to include $\theta$ in the model, which is ignored in the Serkowski formula. Therefore, our model becomes
@@
where $N_{\lambda}$ is the number of different bands for which we had measurements for each target. $q_{i}$, $u_{i}$ is the measurement in a given band, and $\hat{q_{i}}$, $\hat{u_{i}}$ our model introduced in Eq. 6a, 6b. $\sigma_{q_{i}}$, $\sigma_{u_{i}}$ are the uncertainties of the measurements in $q$ and $u$, respectively. The terms $L_{q}$ and $L_{u}$ are the reduced $\chi^{2}$ metric for $q$ and $u$, respectively. The terms $L_{q}/L_{u}$ and $L_{u}/L_{q}$ were introduced to ensure a balanced contribution from $L_{q}$ and $L_{u}$. For example, if the objective was to minimize just $L_{q}+L_{u}$, the optimal value $V$ could be achieved with $L_{q}=V$ and $L_{u}=0$, $L_{u}=V$ and $L_{q}=0$, or any combination in between. By introducing $L_{q}/L_{u}$ and $L_{u}/L_{q}$, we penalize the model when it favors solutions where $L_{q}=0$ or $L_{u}=0$, and thus promote a more balanced fit.
For the modeling procedure, we used the Optuna package in Python (Akiba et al., 2019). Optuna is an automatic hyperparameter optimization software framework, originally developed for machine learning hyperparameter-tuning purposes. However, it can also address general optimization problems through the use of a Bayesian optimization framework. We found it highly effective in multiparameter optimization with few data points and many parameters, such as the case of the $q-u$ modeling, and more efficient in finding the best solution than traditional algorithms such as Markov chain Monte Carlo (MCMC). We used the default options of Optuna with n_trials = 5000 for our fits. Moreover, we adopted the following priors. For $P_{max}$, we initially assumed a rather loose Gaussian prior distribution with a mean of $3\%$ and standard deviation of $3\%$, and enforced it to be positive. For $\lambda_{max}$, we assumed a Gaussian prior distribution with a mean of 0.6 $\mu m$ and standard deviation of 0.2 $\mu m$, and ensured it was always positive. For $K$, we imposed a Gaussian prior distribution with a mean of $1.68\cdot\lambda_{max}$ and standard deviation of $0.13\cdot\lambda_{max}$ (Wilking et al., 1980), and ensured it was always positive. For $\theta$, we assumed a Gaussian prior distribution with the mean the value derived from the average $q$ and $u$ as in Eq. 4 and converted to $\theta$ using Eq. 3b, and a $10^{o}$ standard deviation.
@@ 3.4 Fitting multiple clouds
In order to fit multiple clouds (as in the examples of Sect. 4.2), we again used the $q-u$ fitting technique, following the same logic discussed in Sect. 3.3. The multiple cloud model is derived from Eqs. 2 and 6, by adding up the $q$ and $u$ components of the individual clouds.
When trying to determine the correct parameters for a model with two or more clouds, the number of parameters is increased (four parameters per cloud), resulting in a substantial number of possible combinations and a vast parameter space to explore. In such situations, identifying the optimal combination of parameters that minimize the cost function can be challenging. To address this, we employed the following procedure. Initially, instead of directly using Optuna, we utilized an MCMC algorithm to rapidly explore the parameter space. Given that the two clouds are interchangeable, we imposed the condition $\lambda_{max_{1}}>\lambda_{max_{2}}$ to focus the algorithm on a single solution. Subsequently, we visually examined the corner plots of the parameters generated by the MCMC algorithm to establish reasonable constraints for the Optuna algorithm, thereby significantly narrowing the search space. Finally, we ran the fit using Optuna, imposing Gaussian priors on the parameters, based on the mean and the spread of the parameters identified by the MCMC algorithm. We chose Optuna to derive the final set of parameters for the model, as it directly provides the optimal set of parameters by minimizing Eq. 7. The fits were evaluated with $\chi^{2}_{\nu}$ (Eq. 9).
@@ 4 Results
4.1 Statistical sample
We grouped the sample of sources found in Sect. 3.1 into subsamples depending on the number of clouds ($N_{C}$) lying in their LOS, as derived from the 3D extinction map. We obtained four main subsamples: targets in LOSs with one dominant cloud, $N_{C}=1$; two clouds, $N_{C}=2$; three clouds $N_{C}=3$; and more than three clouds, $N_{C}>3$. For this sample, we fit the traditional Serkowski curve (Eq. 1) for all targets, regardless of $N_{C}$, to evaluate the capability of the Serkowski relation to fit polarimetric data for subsamples with different $N_{C}$. We evaluated the performance of the Serkowski relation in all of the targets as described in Sect. 3.3. We constructed the cumulative density functions (CDFs) of $\chi^{2}_{\nu}$ for the different subsamples (Fig. 5), and we applied the Kolmogorov-Smirnov (KS) test to identify any statistical discrepancies between the $\chi^{2}_{\nu}$ distributions of the $N_{C}=1$ subset and the other three subsets.
Figure 5: CDFs of $\chi^{2}_{\nu}$ for different subsamples, as is described in each label.
We found that the $\chi^{2}_{\nu}$ distribution of the $N_{C}=1$ subset is significantly different from those of subsets with $N_{C}=2$ or $N_{C}=3$. In contrast, the $\chi^{2}_{\nu}$ distributions of the $N_{C}=1$ subset and of the $N_{C}>3$ subset are not discrepant. The $p-$values of the KS tests for all comparisons are quoted in Table 1, along with the median $\chi^{2}_{\nu}$ for the fits of each subsample. Despite the sparse nature of the data (measurements in only $\sim 3.5$ different filters on average per target), the sample is sufficiently large to allow for the statistical distinction of LOSs with two or three clouds from those with only a single cloud. The deterioration of the Serkowski fits in the multicloud LOSs is mainly driven by the behavior of the EVPA with wavelength, and is made clear thanks to our $q-u$ space fits. In contrast, if we perform the fits only on $P-\lambda$, it is not possible to discriminate between LOSs with and without multiple clouds with current data. When the structure in the LOS is too complicated ($N_{C}>3$), there is no statistical difference in the polarization profiles compared to the single cloud case. This is not surprising, as with an increasing number of clouds the variations induced by individual screens tend to average out. Even deviations in the EVPA, which should be more prominent, would become smeared and difficult to detect, especially with limited data points across the spectrum.
Table 1: $p$ values of the KS test between the subsample of targets with $N_{C}=1$ and other subsamples (Sect. 4.1).
Subsamples $p$-value Median $\chi^{2}_{\nu}$
$N_{C}=1$ 1 1.20
$N_{C}=2$ 1.5 $\cdot 10^{-3}$ 2.59
$N_{C}=3$ 2.1 $\cdot 10^{-2}$ 2.17
$N_{C}=2$ or 3 $1.8\cdot 10^{-3}$ 2.51
$N_{C}>1$ 2.1 $\cdot 10^{-2}$ 1.97
$N_{C}>3$ $1.5\cdot 10^{-1}$ 0.86
@@ 4.2 Two revealing examples
To demonstrate the impact of the LOS variations in $\lambda_{max}$ and EVPA on the Serkowski relation, we employed archival spectropolarimetric data (Bagnulo et al., 2017). We identified two revealing cases, where the multiwavelength polarization data show clear evidence of the LOS effects. The examples discussed here are not part of our statistical sample, as their distances are higher than the limitations imposed in Sect. 3.1.
The LOS toward the main sequence star HD93222 with spectral type O7 represents a prominent example of what the $P-\lambda$, and $\theta-\lambda$ profiles look like when 3D effects are significant. First, the polarization as a function of wavelength deviates from the typical Serkowski curve (Fig. 6 middle panel). Fitting the $P-\lambda$ data with the Serkowski relation requires unphysical values for both $\lambda_{max}$, and $K$ (Table 2), which still yield a merely adequate fit. In addition, $\theta$ varies smoothly with wavelength, $\lambda$, which is unexpected when the polarization is dominated by a single cloud with a well-defined mean magnetic field. Thus, both $P$ and EVPA show an atypical behavior with $\lambda$ for this star. Using 3D dust extinction maps (Edenhofer et al., 2024), we verified the existence of multiple clouds along the LOS toward the HD93222 star. The clouds are identified as abrupt increments in extinction as a function of distance (left panel in Fig. 6). The average extinction profile shows three prominent clouds located at around 0.25, 1.35, and 1.8 kpc. The target star, whose estimated distance is 2.44 kpc (Bailer-Jones et al., 2021), lies beyond these clouds, and hence its polarization signal carries information from all of them. We fit the data with models with different number of clouds. Although the extinction profile exhibits three distinct steps, the data can be described equally well with both a two- and a three-cloud model. This may indicate that two out of three clouds have similar properties, and their constructive profile resembles that of one cloud (such as the case of the middle left panel of Fig. 1). We proceeded with the simplest case of a two-cloud model.
Table 2: Fit parameters for the stars HD163181 and HD93222 in Sect. 4.2.
@@
Figure 6: Example of a case in which both the $P-\lambda$ and the $\theta-\lambda$ profiles deviate from Serkowski expectations. We were able to fit both profiles by introducing a 2-Cloud model. Data from Bagnulo et al. (2017) for HD93222. Left: Extinction vs. distance profile taken from the 2kpc version of the map by Edenhofer et al. (2024). The vertical red line corresponds to the distance of the source. Middle: Data in the $P-\lambda$ space, together with three fit models. Solid black line: Fit using Eq. 1 in $P-\lambda$ space only. Dashed blue line: Best-fit 1-Cloud model in the $q-u$ space (Eq. 6). Dash-dotted red line: Best-fit 2-Cloud model in the $q-u$ space (Eq. 6). Dotted orange lines correspond to the two individual components of the 2-Cloud model. Right: Data in $\theta-\lambda$ space together with the three models and the individual components of the 2-Cloud model. Parameters of the fits are quoted in Table 2.
@@
As a second example, we show results for the star HD163181, which is an intermediate-size luminous supergiant of spectral class O9.5 (Fig. 7). In this case, the observed degree of polarization follows the Serkowski formula, but the EVPA varies with wavelength. The 3D extinction profile has two significant steps, indicating the existence of more than one dominant cloud along this LOS. Assuming the standard Serkowski curve with a single dominant cloud can yield good fits to the data in the $P-\lambda$ space (Bagnulo et al., 2017), but it fails to explain the observed variability of the EVPA; we verified this by fitting a one cloud model in the $q-u$ plane (dashed blue line in Fig. 7). However, when we considered two clouds along this LOS, as the 3D extinction map suggests, we were able to capture the observed variability of both $P$ and $\theta$ with wavelength (dash-dotted red line in Fig. 7, middle panel).
@@
Figure 7: Example of a case in which the $P-\lambda$ profile follows the Serkowski relation, but the EVPA does not. By employing a model with two components in the LOS, it is possible to fit both $P$ and $\theta$. Data from Bagnulo et al. (2017) for HD163181. Panels and lines similar to Fig. 6. Parameters of the fits are quoted in Table 2.
4.3 The parameters of the Serkowski formula